OpenStudy (anonymous):

Given f(x)=3x+1 and (f+g)(x)=6-1/2x, find the function of g.

OpenStudy (jamesj):

g(x) = (f+g)(x) - f(x) = 6 - 1/2x - (3x + 1) = ....